# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ I Introduction
With the discovery of the Higgs boson at the CERN Large Hadron Collider (LHC) about ten years ago [1, 2], the full particle content of the Standard Model (SM) of particle physics has become fully fixed. Among the 17 existing elementary particles (6 quarks, 6 leptons, 4 gauge bosons, and the scalar boson), the top quark, the Higgs and the electroweak (W, Z) bosons are the most massive ones. Studying in detail the properties of the four heaviest elementary particles, with masses around the electroweak scale $\Lambda_{\mathrm{EW}}\approx\mathcal{O}(100$ GeV), is an important priority in precision SM studies and in searches for new physics beyond it (BSM). At the LHC, the large center-of-mass energies (up to $\sqrt{s}=14$ TeV) and integrated luminosities (up to $\mathcal{L}_{\mathrm{int}}=3$ ab-1 at the end of the high-luminosity, HL-LHC, phase) [3] available in proton-proton (p-p) collisions, as well as the many ab-1 to be integrated in the very clean “background-free” $\mathrm{e}^{+}\mathrm{e}^{-}$ collision environment of the next planned lepton collider facilities, such as the Future Circular Collider (FCC-ee) [4] or the CEPC [5], will allow the collection of very large H, W, Z, and top-quark data samples. In addition, and despite much more complicated background conditions than at $\mathrm{e}^{+}\mathrm{e}^{-}$ colliders, p-p collisions at $\sqrt{s}=100$ TeV planned at FCC-hh [6] will produce unprecedented numbers of W, Z, Higgs, and top particles. The very large data samples expected at these machines will make it possible to measure many of their rare few-body decays —understood here as decays into two to four final-state particles, with branching fractions below $\mathcal{B}\approx 10^{-5}$—, which remain unobserved to date. Broadly speaking, in this work we consider three types of rare few-body decays shown in Fig. 1: (i) decays into three (or four) lighter gauge bosons, or into one (or two) gauge boson plus two neutrinos, (ii) decays into a lighter gauge boson plus a single hadronic (or leptonic) bound system in the form of a quarkonium (or leptonium) state, and (iii) exclusive decays into two quarkonium bound states. The reason for the rarity of the first type of decays is the fact that they proceed through suppressed (heavy particle) loops, whereas the second and third decay modes occur very scarcely because the probability to form a single (let alone double) -onium bound state, out of the outgoing quarks or leptons of the primary decay, is very small.
@@
The incentives for the study of such decays are multiple, with some specificities depending on the decaying “mother” particle. A first general motivation for their study is the possibility that new physics phenomena alter some very rare partial decay widths. Precision tests of suppressed (or forbidden) processes in the SM —such as flavour changing neutral currents (FCNC), or processes violating lepton flavour (LFV) or lepton flavour universality (LFUV)— are powerful probes of BSM physics that have been mostly studied so far in b-quark decays [7, 8]. Unlike in the latter case where, due to the relatively low masses of the B hadrons involved, large power corrections to the decay rates lead to sizable theoretical uncertainties, power corrections in decays of the electroweak and Higgs bosons and top quark are under better theoretical control thanks to the large boost of the final-state hadrons. More concrete motivations are succinctly mentioned next for each particle. In the case of the Higgs boson, very rare decays with photons and/or neutrinos (Fig. 1, left) lead to experimental signatures that are identical to potential exotic BSM Higgs decays [9, 10], and therefore need to be well estimated as background(s) for the latter. Exclusive Higgs decays such as those shown in Fig. 1 (center) are sensitive to the Yukawa couplings of the charm and lighter quarks [11, 12, 13, 14, 15], as well as to FCNC $\rm H\to qq^{\prime}$ decays, which are otherwise very difficult to access experimentally due to the smallness of the quark masses involved, and/or their heavily suppressed loop-induced rates [10]. In the Z boson case, specific decay modes allow probing FCNC couplings (which are loop- and Glashow–Iliopoulos–Maiani (GIM) [16] suppressed processes in the SM) in a model-independent way [17, 18]. In addition, rare decays of the H and Z bosons shown in Fig. 1 where the onium state decays into diphotons or dileptons, constitute a background for different searches for exotic BSM decays [19], such as e.g., $\mathrm{Z}\to\gamma\,a(\gamma\gamma)$ with $a$ being an axion-like particle [20] or a massive graviton [21] decaying into photons; or $\rm H\to A^{\prime}(\ell^{+}\ell^{-})+X$ where a dark photon A further decays into a $\ell^{+}\ell^{-}$ lepton pair [22]. The measurements of exclusive decays of the W boson [23, 24, 25] and of the top quark [26] have also been suggested e.g., as alternative means to determine the W boson and top quark masses via two- or three-body invariant mass analysis, free of invisible neutrinos or (messy) jets involved in the inclusive decay modes. In the top quark case, the study of its suppressed radiative decay rates, induced by the offdiagonal parts of the quark dipole moments, has been of interest for many decades because they also provide an experimentally clean probe for new physics [27]. Of particular importance are precision studies of rare FCNC decays such as $\rm t\to Zc$, $\rm t\to\gamma c$, and $\rm t\to cg$, for which many BSM extensions can enhance their branching ratios by orders of magnitude, thereby yielding compelling phenomenology [27, 28].
@@
Theoretically, the calculation of the partial widths of the few-body decays schematically shown in Fig. 1 (left) are carried out through an expansion of the underlying virtual loops in the electroweak (EW) and/or quantum chromodynamics (QCD) couplings ($\alpha$ and $\alpha_{\text{s}}$, respectively), at a given order of perturbative accuracy. First calculations were computed at leading order (LO), but more recent results exist at next-to-leading-order (NLO) accuracy. With regards to the center and right diagrams of Fig. 1, the formalism of QCD factorization [29, 30, 31, 32, 33] is a well-established approach to study and compute rates for hard exclusive processes with individual hadrons in the final state. Within this framework, the production of the final state occurs in two stages: first, a pair of quarks is produced in a short-distance partonic process; second, the quarks form a given hadronic bound state. The short-distance physics at the energy scale of the initial heavy particle is appropriately separated from the long-distance dynamics governing the formation of the final hadron(s), and the decay amplitudes can be therefore obtained from the convolution of hard-scattering functions calculable in perturbative QCD (pQCD) and nonlocal hadronic matrix elements. These latter objects, called light-cone distribution amplitudes (LCDAs), are nonperturbative and scale-dependent functions that encode the infrared physics of the final hadron formation. The decay amplitudes are formally given as expansions in the ratio of the two (hard and soft) scales in the problem, given by the large energy released $E=m_{\mathrm{X}}/2$ in the process and the final hadron mass, respectively. In the cases of interest in this work, the hard scale is set by the heavy mass of the decaying bosons or top quark, $\mu_{\text{hard}}\approx\Lambda_{\mathrm{EW}}$ rendering the impact of nonperturbative power corrections at typical hadronization energy scales $\Lambda_{\text{QCD}}\approx 0.2$ GeV, $\mathcal{O}(\Lambda_{\text{QCD}}/\Lambda_{\mathrm{EW}})\lesssim 10^{-3}$, under control. The studies of exclusive decays of the H, Z, W bosons, and top quark, therefore not only provide a sensitive test of the SM but, in particular, also stringent tests of the QCD factorization formalism, including constraints on poorly known aspects of the nonperturbative formation of hadronic bound states. In the case of final states with charm and bottom quarks, they can bring forward valuable insights into partially conflicting mechanisms of heavy quarkonium production [34, 35].
@@ II.1.1 Light-cone (LC) factorization
A common method for computing exclusive light-hadron production in high-energy decays is given by the so-called amplitude expansion in the light cone (LC), where the perturbatively small expansion parameter is the chirality factor $m_{\mathrm{q}}/E$ (with $m_{\mathrm{q}}$ and $E$ being the mass of the produced quark and its typical energy, respectively), and the meson form-factors are nonperturbative objects described by LCDAs [33] that encode the internal motion of the quark-antiquark pair inside their bound state, constrained by QCD sum rules [36, 37]. The LC approach implements directly QCD factorization by describing the production of a given final state as a convolution of a hard scattering kernel times the LCDA object. As an example, for the case of the exclusive decay of a heavy particle X into a meson M plus a photon ($\rm X\to M+\gamma$), the decay amplitude can be written schematically as
@@
[…]
@@
[…]
@@
Higher-order pQCD corrections can be included in Eq. (1) through the evolution of $H_{\mathrm{M}}(x,\mu)$ from $m_{\mathrm{X}}$ down to the meson mass scale. In the simplest approximation where one ignores the internal motion of quarks inside the mesons, the LCDA takes the form $\phi(x,\mu)=\delta(x)$ or $\phi(x,\mu)=x\delta(x)$ for parity-even and odd states, respectively, where $\delta(x)$ is the Dirac delta function, and the whole meson information in Eq. (1) is encoded in the $f_{\mathrm{M}}$ parameter. In such a delta-approximation scheme, one can then write e.g., the partial decay width of a Z boson into an exclusive $\mathrm{J}/\psi+\gamma$ final state as
@@ II.1.2 Soft-Collinear Effective Theory (SCET)
As an alternative to the LC approach, many decay modes involving light mesons have been calculated using Soft-Collinear Effective Theory (SCET) [38], where QCD factorization is rephrased in the language of effective field theory (EFT) [39] to properly address the problem of the multiple scales appearing in the calculations. The SCET framework provides a systematic expansion of $\rm X\to M+Y$ decay amplitudes in powers of a small expansion parameter $\lambda=\Lambda_{\text{QCD}}/E$ (with $E=m_{\mathrm{X}}/2$ the meson energy), and allows proper resummation of the large logarithms appearing in the ratios of scales $\log(m_{\mathrm{M}}/m_{\mathrm{X}})^{n}$, given by the mass of the mother particle and that of the mesonic final state, which spoil the convergence of the perturbative expansion. Most often, SCET is combined with LCDA. At the scale of the large energies released in EW and Higgs boson decays, even charm and bottom quarks can be treated as light quarks, and hence heavy-quark mesons can be described by LCDAs. In a few other calculations, the EFT approach is combined with form-factors based on the nonrelativistic quark model (NRQM) [40]. Final-state mesons containing one light and one heavy quark can be instead described in Heavy-Quark Effective Theory (HQET) where the LCDA describes the hadronic physics at two distinct scales, $m_{\mathrm{Q}}$ and $\Lambda_{\text{QCD}}$, of which the former should be tractable by perturbative methods [41]. In HQET, the hadronic matrix elements are expressed as a combination of perturbatively computable coefficients and new, suitably defined, hadronic matrix elements that exploit the constraints provided by heavy quark symmetries [42] on the nonperturbative matrix elements at low scales.
At leading order in the SCET expansion, the $\rm X\to M+\gamma$ decay amplitude can be written in a factorized form as [43]:
@@
where $i$ runs over different combinations of gamma matrices $\Gamma_{i}$ that have the appropriate spin structure to form the meson M with wavefunction $M(k)$ from a suitable quark-antiquark pair $(\mathrm{q}\overline{\mathrm{q}})$ combination; $n,\bar{n}$ are four-momenta in the lightcone basis [^1] In the rest frame of a decaying particle X (with mass $m_{\mathrm{X}}\gg m_{\mathrm{M}}$), the 4-momenta of the two final-state particles are $k_{1}^{\mu}=n^{\mu}E$ and $k_{2}^{\mu}=\bar{n}^{\mu}E$, where $E=m_{\mathrm{X}}/2$ is the energy of the final particles in the X rest frame, and the light-cone vectors satisfy $n\cdot\bar{n}=2$.; the Wilson coefficients $C_{i}(t,\mu)$ are calculable hard-scattering EFT coefficients that depend on the process and are evaluated at a factorization scale $\mu$ that can be determined in a matching computation from perturbation theory. The meson matrix element in Eq. (5) has a direct connection with the meson form factor given by Eqs. (2), through [44]:
$\langle M(k)|\bar{q}(t\bar{n})\frac{\bar{n}\mkern-7.5mu/}{2}\Gamma_{i}[t\bar{n},0]q(0)|0\rangle$ $=-if_{\mathrm{M}}E\int_{0}^{1}\mathrm{d}xe^{ixt\bar{n}\cdot k}\phi_{\mathrm{M}}(x,\mu);\quad\rm M=P,VM_{||}$ (6)
@@
where $E=\bar{n}\cdot k/2$; and ,|$\perp$,||⟂ , | | stand for transverse and longitudinal polarizations of the meson. These last expressions can be thought of as the generalization of the relation of matrix elements to decay constants in the nonlocal case: in the local limit $t\to 0$, the matrix elements of Eq. (6) translate into Eq. (2). To make the connection to the LC approach clearer, we can further take the Fourier transform of $C_{i}(t,\mu)$ and call it hard scattering kernel $H_{\mathrm{M}}(x,\mu)\equiv\int\mathrm{d}t\,C_{\mathrm{M}}(t,\mu)\,e^{ixt\bar{n}\cdot k}$. We then have a form of the factorized scattering amplitude
@@
which looks like the LC decay amplitude given by the convolution of hard-scattering kernels with meson LCDA of Eq. (1). One can also expand the LCDAs in terms of Gegenbauer polynomials as [44]
@@
with coefficients $a_{n}^{\mathrm{M}}(\mu)$, moments of the LCDA, that can be obtained from nonperturbative approaches such as lattice QCD, or QCD sum rules. At the electroweak scale, the LCDAs are close to the asymptotic $6x(1-x)$ form. Thus for the asymptotic case, one has e.g., that the $\rm Z\to\mathrm{J}/\psi+\gamma$ partial decay width amounts to [43]:

### v4

@@ I Introduction
With the discovery of the Higgs boson at the CERN Large Hadron Collider (LHC) about ten years ago [ATLAS:2012yve, CMS:2012qbp], the full particle content of the Standard Model (SM) of particle physics has become fully fixed. Among the 17 existing elementary particles (6 quarks, 6 leptons, 4 gauge bosons, and the scalar boson), the top quark, the Higgs and the electroweak (W, Z) bosons are the most massive ones. Studying in detail the properties of the four heaviest elementary particles, with masses around the electroweak scale $\Lambda_{\mathrm{EW}}\approx\mathcal{O}(100$ GeV), is an important priority in precision SM studies and in searches for new physics beyond it (BSM). At the LHC, the large center-of-mass energies (up to $\sqrt{s}=14$ TeV) and integrated luminosities (up to $\mathcal{L}_{\mathrm{int}}=3$ ab-1 at the end of the high-luminosity, HL-LHC, phase) [Azzi:2019yne] available in proton-proton (p-p) collisions, as well as the many ab-1 to be integrated in the very clean “background-free” $\mathrm{e}^{+}\mathrm{e}^{-}$ collision environment of the next planned lepton collider facilities, such as the Future Circular Collider (FCC-ee) [FCC:2018evy] or the CEPC [CEPCStudyGroup:2018ghi], will allow the collection of very large H, W, Z, and top-quark data samples. In addition, and despite much more complicated background conditions than at $\mathrm{e}^{+}\mathrm{e}^{-}$ colliders, p-p collisions at $\sqrt{s}=100$ TeV planned at FCC-hh [FCC:2018vvp] will produce unprecedented numbers of W, Z, Higgs, and top particles. The very large data samples expected at these machines will make it possible to measure many of their rare few-body decays —understood here as decays into two to four final-state particles, with branching fractions below $\mathcal{B}\approx 10^{-5}$—, which remain unobserved to date. Broadly speaking, in this work we consider three types of rare few-body decays shown in Fig. 1: (i) decays into three (or four) lighter gauge bosons, or into one (or two) gauge boson plus two neutrinos, (ii) decays into a lighter gauge boson plus a single hadronic (or leptonic) bound system in the form of a quarkonium (or leptonium) state, and (iii) exclusive decays into two quarkonium bound states. The reason for the rarity of the first type of decays is the fact that they proceed through suppressed (heavy particle) loops, whereas the second and third decay modes occur very scarcely because the probability to form a single (let alone double) -onium bound state, out of the outgoing quarks or leptons of the primary decay, is very small.
@@
The incentives for the study of such decays are multiple, with some specificities depending on the decaying “mother” particle. A first general motivation for their study is the possibility that new physics phenomena alter some very rare partial decay widths. Precision tests of suppressed (or forbidden) processes in the SM —such as flavour changing neutral currents (FCNC), or processes violating lepton flavour (LFV) or lepton flavour universality (LFUV)— are powerful probes of BSM physics. While competitive studies exist at low energies (e.g., in rare kaon decays, LFV searches with muons or taus, or electric dipole moment measurements), most investigations at multi-GeV mass scales have so far been mostly carried out exploiting b-quark decays at colliders [LHCb:2018roe, Belle-II:2018jsg]. Unlike in the latter case where, due to the relatively low masses of the B hadrons involved, large power corrections to the decay rates lead to sizable theoretical uncertainties, power corrections in decays of the electroweak and Higgs bosons and top quark are under better theoretical control thanks to the large boost of the final-state hadrons. More concrete motivations are succinctly mentioned next for each particle. In the case of the Higgs boson, very rare decays with photons and/or neutrinos (Fig. 1, left) lead to experimental signatures that are identical to potential exotic BSM Higgs decays [Curtin:2013fra, LHCHiggsCrossSectionWorkingGroup:2016ypw], and therefore need to be well estimated as background(s) for the latter. Exclusive Higgs decays such as those shown in Fig. 1 (center) are sensitive to the Yukawa couplings of the charm and lighter quarks [Bodwin:2013gca, Isidori:2013cla, Kagan:2014ila, Konig:2015qat, Perez:2015lra], as well as to FCNC $\rm H\to qq^{\prime}$ decays, which are otherwise very difficult to access experimentally due to the smallness of the quark masses involved, and/or their heavily suppressed loop-induced rates [LHCHiggsCrossSectionWorkingGroup:2016ypw]. In the Z boson case, specific decay modes allow probing FCNC couplings (which are loop- and Glashow–Iliopoulos–Maiani (GIM) [Glashow:1970gm] suppressed processes in the SM) in a model-independent way [Guberina:1980dc, Perez:2003ad]. In addition, rare decays of the H and Z bosons shown in Fig. 1 where the onium state decays into diphotons or dileptons, constitute a background for different searches for exotic BSM decays [Agrawal:2021dbo], such as e.g., $\mathrm{Z}\to\gamma\,a(\gamma\gamma)$ with $a$ being an axion-like particle [dEnterria:2021ljz] or a massive graviton [dEnterria:2023npy] decaying into photons; or $\rm H\to A^{\prime}(\ell^{+}\ell^{-})+X$ where a dark photon A further decays into a $\ell^{+}\ell^{-}$ lepton pair [Curtin:2014cca]. The measurements of exclusive decays of the W boson [Arnellos:1981gy, Mangano:2014xta, Jones:2020bvu] and of the top quark [dEnterria:2020ygk] have also been suggested e.g., as alternative means to determine the W boson and top quark masses and widths via two- or three-body invariant mass analyses, free of invisible neutrinos or (messy) jets involved in the inclusive decay modes. In the top quark case, the study of its suppressed radiative decay rates, induced by the offdiagonal parts of the quark dipole moments, has been of interest for many decades because they also provide an experimentally clean probe for new physics [Beneke:2000hk]. Of particular importance are precision studies of rare FCNC decays such as $\rm t\to Zc$, $\rm t\to\gamma c$, and $\rm t\to cg$, for which many BSM extensions can enhance their branching ratios by orders of magnitude, thereby yielding compelling phenomenology [Beneke:2000hk, Aguilar-Saavedra:2004mfd].
@@
Theoretically, the calculation of the partial widths of the few-body decays schematically shown in Fig. 1 (left) are carried out through an expansion of the underlying virtual loops in the electroweak (EW) and/or quantum chromodynamics (QCD) couplings ($\alpha$ and $\alpha_{\text{s}}$, respectively), at a given order of perturbative accuracy. First calculations were computed at leading order (LO), but more recent results exist at next-to-leading-order (NLO) accuracy. With regards to the center and right diagrams of Fig. 1, the formalism of QCD factorization [Lepage:1979zb, Lepage:1980fj, Efremov:1978rn, Efremov:1979qk, Chernyak:1983ej] is a well-established approach to study and compute rates for hard exclusive processes with individual hadrons in the final state. Within this framework, the production of the final state occurs in two stages: first, a pair of quarks is produced in a short-distance partonic process; second, the quarks form a given hadronic bound state. The short-distance physics at the energy scale of the initial heavy particle is appropriately separated from the long-distance dynamics governing the formation of the final hadron(s), and the decay amplitudes can be therefore obtained from the convolution of hard-scattering functions calculable in perturbative QCD (pQCD) and nonlocal hadronic matrix elements. These latter objects, called light-cone distribution amplitudes (LCDAs), are nonperturbative and scale-dependent functions that encode the infrared physics of the final hadron formation. The decay amplitudes are formally given as expansions in the ratio of the two (hard and soft) scales in the problem, given by the large energy released $E=m_{\mathrm{X}}/2$ in the process and the final hadron mass, respectively. In the cases of interest in this work, the hard scale is set by the heavy mass of the decaying bosons or top quark, $\mu_{\text{hard}}\approx\Lambda_{\mathrm{EW}}$ rendering the impact of nonperturbative power corrections at typical hadronization energy scales $\Lambda_{\text{QCD}}\approx 0.2$ GeV, $\mathcal{O}(\Lambda_{\text{QCD}}/\Lambda_{\mathrm{EW}})\lesssim 10^{-3}$, under control. The studies of exclusive decays of the H, Z, W bosons, and top quark, therefore not only provide a sensitive test of the SM but, in particular, also stringent tests of the QCD factorization formalism, including constraints on poorly known aspects of the nonperturbative formation of hadronic bound states. In the case of final states with charm and bottom quarks, they can bring forward valuable insights into partially conflicting mechanisms of heavy quarkonium production [Lansberg:2019adr, Chapon:2020heu].
@@ II.1.1 Light-cone (LC) factorization
A common method for computing exclusive light-hadron production in high-energy decays is given by the so-called amplitude expansion in the light cone (LC), where the perturbatively small expansion parameter is the chirality factor $m_{\mathrm{q}}/E$ (with $m_{\mathrm{q}}$ and $E$ being the mass of the produced quark and its typical energy, respectively), and the meson form-factors are nonperturbative objects described by LCDAs [Chernyak:1983ej] that encode the internal motion of the quark-antiquark pair inside their bound state, constrained by QCD sum rules [Shifman:1978bx, Colangelo:2000dp]. The LC approach implements directly QCD factorization by describing the production of a given final state as a convolution of a hard scattering kernel times the LCDA object. As an example, for the case of the exclusive decay of a heavy particle X into a meson M plus a photon ($\rm X\to M+\gamma$), the decay amplitude can be written schematically as
@@
where $f_{\mathrm{M}}$ is the meson decay constant, $x$ is the difference in the momentum fractions carried out by the quark and the antiquark inside the outgoing meson, $H_{\mathrm{M}}(x,\mu)$ is the hard scattering kernel, calculable in pQCD from the relevant matrix-element diagrams, and evaluated at the energy scale $\mu$ of the decay, and $\phi_{\mathrm{M}}(x)$ is the LCDA object, which at LO can be interpreted as the amplitude for finding a quark-antiquark pair with longitudinal momentum fraction difference $x$ inside the meson. The $f_{\mathrm{M}}$ constant is a parameter that describes the strength of the interaction between a meson and its decay products. It is connected with the form factor of the SM transition current sandwiched between the vacuum and the meson states, which for pseudoscalar (P) and vector meson (VM) mesons (with masses $m_{\mathrm{P}},m_{\mathrm{VM}}$ and wavefunctions $P(k)$ and $\mathit{VM}(k)$ at four-momentum $k$) read, respectively,
@@
$\langle\mathit{VM}(k)|{\overline{q}}_{1}(0)\gamma^{\mu}q_{2}(0)|0\rangle$ $=-if_{\mathrm{VM}}m_{\mathrm{VM}}\varepsilon^{*\mu}_{\mathrm{VM}}.$
@@
Higher-order pQCD corrections can be included in Eq. (1) through the evolution of $H_{\mathrm{M}}(x,\mu)$ from $m_{\mathrm{X}}$ down to the meson mass scale. In the simplest approximation where one ignores the internal motion of quarks inside the mesons, the LCDA takes the form $\phi(x,\mu)=\delta(x)$ or $\phi(x,\mu)=x\delta(x)$ for parity-even and odd states, respectively, where $\delta(x)$ is the Dirac delta function, and the whole meson information in Eq. (1) is encoded in the $f_{\mathrm{M}}$ parameter. In such a delta-approximation scheme, one can then write e.g., the partial decay width of a Z boson into an exclusive $\mathrm{J}/\psi+\gamma$ final state as [Luchinsky:2017jab]
@@ II.1.2 Soft-Collinear Effective Theory (SCET)
As an alternative to the LC approach, many decay modes involving light mesons have been calculated using Soft-Collinear Effective Theory (SCET) [Bauer:2001yt], where QCD factorization is rephrased in the language of effective field theory (EFT) [Georgi:1993mps] to properly address the problem of the multiple scales appearing in the calculations. The SCET framework provides a systematic expansion of $\rm X\to M+Y$ decay amplitudes in powers of a small expansion parameter $\lambda=\Lambda_{\text{QCD}}/E$ (with $E=m_{\mathrm{X}}/2$ the meson energy), and allows proper resummation of the large logarithms appearing in the ratios of scales $\log^{n}(m_{\mathrm{M}}/m_{\mathrm{X}})$, given by the mass of the mother particle and that of the mesonic final state, which spoil the convergence of the perturbative expansion. Most often, SCET is combined with LCDA. At the scale of the large energies released in EW and Higgs boson decays, even charm and bottom quarks can be treated as light quarks, and hence heavy-quark mesons can be described by LCDAs. In a few other calculations, the EFT approach is combined with form-factors based on the nonrelativistic quark model (NRQM) [Manohar:1983md]. Final-state mesons containing one light and one heavy quark can be instead described in Heavy-Quark Effective Theory (HQET) where the LCDA describes the hadronic physics at two distinct scales, $m_{\mathrm{Q}}$ and $\Lambda_{\text{QCD}}$, of which the former should be tractable by perturbative methods [Beneke:2000ry]. In HQET, the hadronic matrix elements are expressed as a combination of perturbatively computable coefficients and new, suitably defined, hadronic matrix elements that exploit the constraints provided by heavy quark symmetries [Neubert:1993mb] on the nonperturbative matrix elements at low scales.
At leading order in the SCET expansion, the $\rm X\to M+\gamma$ decay amplitude can be written in a factorized form as [Konig:2018wuf]:
@@
where $i$ runs over different combinations of gamma matrices $\Gamma_{i}$ that have the appropriate spin structure to form the meson M with wavefunction $M(k)$ from a suitable quark-antiquark pair $(\mathrm{q}\overline{\mathrm{q}})$ combination; $n,\bar{n}$ are four-momenta in the lightcone basis [^1] In the rest frame of a decaying particle X (with mass $m_{\mathrm{X}}\gg m_{\mathrm{M}}$), the 4-momenta of the two final-state particles are $k_{1}^{\mu}=n^{\mu}E$ and $k_{2}^{\mu}=\bar{n}^{\mu}E$, where $E=m_{\mathrm{X}}/2$ is the energy of the final particles in the X rest frame, and the light-cone vectors satisfy $n\cdot\bar{n}=2$.; the Wilson coefficients $C_{i}(t,\mu)$ are calculable hard-scattering EFT coefficients that depend on the process and are evaluated at a factorization scale $\mu$ that can be determined in a matching computation from perturbation theory. The meson matrix element in Eq. (5) has a direct connection with the meson form factor given by Eqs. (2), through [Grossman:2015cak]:
@@
$\langle\mathit{VM}(k)_{\perp}|\bar{q}(t\bar{n})\frac{\bar{n}\mkern-7.5mu/}{2}\Gamma_{i}[t\bar{n},0]q(0)|0\rangle$ $=-if^{\perp}_{\mathrm{VM}}(\mu)E\varepsilon_{\mathrm{VM}}^{\perp*\mu}\int_{0}^{1}\mathrm{d}xe^{ixt\bar{n}\cdot k}\phi^{\perp}_{\mathrm{VM}}(x,\mu),$
where $E=\bar{n}\cdot k/2$; and $\perp,||$ stand for transverse and longitudinal polarizations of the meson. These last expressions can be thought of as the generalization of the relation of matrix elements to decay constants in the nonlocal case: in the local limit $t\to 0$, the matrix elements of Eq. (6) translate into Eq. (2). To make the connection to the LC approach clearer, we can further take the Fourier transform of $C_{i}(t,\mu)$ and call it hard scattering kernel $H_{\mathrm{M}}(x,\mu)\equiv\int\mathrm{d}t\,C_{\mathrm{M}}(t,\mu)\,e^{ixt\bar{n}\cdot k}$. We then have a form of the factorized scattering amplitude
@@
which looks like the LC decay amplitude given by the convolution of hard-scattering kernels with meson LCDA of Eq. (1). One can also expand the LCDAs in terms of Gegenbauer polynomials as [Grossman:2015cak]
@@
with coefficients $a_{n}^{\mathrm{M}}(\mu)$, moments of the LCDA, that can be obtained from nonperturbative approaches such as lattice QCD, or QCD sum rules. At the electroweak scale, the LCDAs are close to the asymptotic 6x(1x)6x(1-x) form. Thus for the asymptotic case, one has e.g., that the $\rm Z\to\mathrm{J}/\psi+\gamma$ partial decay width amounts to [Konig:2018wuf]: